OpenStudy (anonymous):

Im really confused on how to factor this one 64t^6-27v^3 I need step by step

OpenStudy (imstuck):

That is supposed to be two different variables in there? A t and a v?

OpenStudy (anonymous):

yeah

OpenStudy (skullpatrol):

It looks like a difference of cubes: $$\Huge x^3-y^3=(x-y)(x^2+xy+y^2)$$

OpenStudy (skullpatrol):

$$\Huge 64t^6-27v^3=(4t^2)^3-(3v)^3$$

OpenStudy (anonymous):

how do I factor (a-b)^3+27

OpenStudy (skullpatrol):

That is the sum of two cubes: $$x^3+y^3=(x+y)(x^2-xy+y^2)$$

OpenStudy (anonymous):

you have to factor it using the difference of two cubes

OpenStudy (anonymous):

I'll write the answer if that helps how to factor it

OpenStudy (anonymous):

this is the answer (a-b+3)(a^2-2ab+b^2-3a+3b+9)